OpenStudy (anonymous):

Solve the congruence: 19x + 7 congreunce 27 (mod 18).

jimthompson5910 (jim_thompson5910):

19x + 7 = 27 (mod 18) 1x + 7 = 9 (mod 18) 1x = 9-7 (mod 18) 1x = 2 (mod 18) x = 2 (mod 18)

OpenStudy (anonymous):

How did you get from step one to two

jimthompson5910 (jim_thompson5910):

19 = 1 (mod 18) 27 = 9 (mod 18)

jimthompson5910 (jim_thompson5910):

19 = 1 (mod 18) because 19/18 = 1 remainder 1 ------------------------------------------------------- 27 = 9 (mod 18) because 27/18 = 1 remainder 9

jimthompson5910 (jim_thompson5910):

so that's how 19x + 7 = 27 (mod 18) turned into 1x + 7 = 9 (mod 18)

OpenStudy (anonymous):

Ok got it